The shelf life of a particular dairy product is normally distributed with a mean of 12 days and a standard deviation of 3 days. a) About what percent of the products last between 10 and 12 days


1
Expert's answer
2021-10-27T15:23:59-0400

The shelf life has the next distribution: "N(12,3^{2})=12+3N(0,1)"

a) About what percent of the products last between 10 and 12 days

The point is to find "P(10<12+3N(0,1)<12)"

"P(10<12+3N(0,1)<12) = P(-{\frac 2 3}<N(0,1)<0)=P(N(0,1)<0)-P(N(0,1)<-{\frac 2 3}) = 0.5 - 0.2525=0.2475"

About 24.75 percents of products will last between 10 and 12 days